# On the Trade-Off between Stability and Fidelity of Gaussian-Smoothed Saliency Maps

### v1

@@ 4.3 Fidelity of Gaussian Smoothed Saliency Maps
We have demonstrated that increasing $\sigma$ enhances the algorithmic stability of Smooth-Grad. However, in this subsection, we illustrate that this stability improvement comes at the expense of fidelity. As $\sigma$ increases, the Gaussian-smoothed saliency map diverges further from the ground-truth map.
First, we formally define the fidelity error of the $\sigma$-smoothed saliency map algorithm $\operatorname{Sal}$ as the largest possible expected $\ell_{2}$-distance between the smoothed and unsmoothed saliency maps for any training set $S$, labeled data sample $x,y$, where the expectation is over the randomness of the algorithm.
$\epsilon_{\mathrm{fidelity}}(\operatorname{Sal},\sigma):=\sup_{S,x,y}\mathbb{E}_{A}\Bigl{[}\Big{\|}$ $\mathbb{E}_{z\sim N(0,\sigma^{2}I)}\left[\operatorname{Sal}_{A(S)}(x+z,y)\right]-\operatorname{Sal}_{A(S)}(x,y)\Big{\|}\Bigr{]}$ (11)
@@ 5 Numerical Experiments
Figure 2: Relationship between stability error of saliency maps and sigma, evaluated on the ImageNet test set. The shaded region indicates one standard deviation from the mean value. Here $x_{\text{max}}-x_{\text{min}}$ represents the largest value range for all pixels, following [31].
@@
The goal of our numerical experiments is to validate the stability-fidelity tradeoff incurred by choosing different smoothing factors $\sigma$. We evaluated different neural network architectures and different training set sizes on CIFAR10 [18] and ImageNet [7] datasets. We trained ResNet34 [14] for CIFAR10 and ResNet50 for ImageNet, due to different dataset scales. More detailed experiment configurations are in Appendix C.
@@ 5.1 Saliency Map Stability and Fidelity Numerical Results
$\mathbb{E}_{(x,y)\sim D}\mathbb{E}_{i=1}^{2}\bigl{\|}$ $\mathbb{E}_{z\sim N(0,\sigma^{2}I)}[\operatorname{Sal}_{A(S_{i})}(x+z,y)]-\operatorname{Sal}_{A(S_{i})}(x,y)\bigr{\|}$ (13)
@@ 5.2 Visualizations of Saliency Map Stability and Fidelity
Additionally, we visualize the trend that a larger $\sigma$ results in a smoothed saliency map that is less similar to the vanilla version, as demonstrated in Figure 15 for both Simple-Grad and Integrated-Grad. The visualization results, along with our theoretical findings, indicate that Gaussian smoothing enhances the stability of saliency maps at the expense of fidelity. Therefore, the choice of $\sigma$ in practical applications should be carefully determined by considering this trade-off.
@@ 5.3 Empirical Results on Generalized Settings
Beyond the randomness of training settings, we further examine other sources that could cause instability in interpretation: different training recipes and model architectures. Extensive experiments across various settings, as shown in Table 1, empirically indicate that the stabilizing effect of Gaussian smoothing also holds in different scenarios. We use the structural similarity index measure SSIM [34] to evaluate the similarity of two saliency maps and a larger SSIM value indicates that the two maps are more perceptually similar. We additionally use top-k mIoU to evaluate the similarity by calculating mIoU between the most salient $k=500$ pixels of two maps.
In the top four rows, we train two ResNet50 models on two disjoint ImageNet training subsets of 500,000 samples each. For all other rows, we use pre-trained checkpoints provided by TorchVision [24]. Since TorchVision provides two ResNet50 checkpoints trained with different schemes, we use this pair for the middle four rows of our table. For the ResNet50 models in the bottom rows, we use the checkpoint version 1.
First, when networks are trained on different training sets, Gaussian smoothing makes the resulting saliency maps more similar for both Simple-Grad and Integrated-Grad. For advanced saliency map algorithms such as Grad-CAM [28] and Grad-CAM++ [6], the vanilla maps already have high SSIM values thus are quite similar. Interestingly, applying Gaussian smoothing occasionally causes significant visual changes in one saliency map but not the other, resulting in a slightly lower SSIM value for Grad-CAM. Second, the instability caused by different training recipes can be largely mitigated by Gaussian smoothing, with better SSIM and top-k mIoU scores. See the corresponding visualization for Simple-Grad in Figure 8 in Appendix A.
Third, the effect of smoothing holds for most cross-architecture experiments. Besides ResNet, we also evaluated more recent architectures such as Swin Transformer [22] and ConvNeXt [23]. See Figures 9, 10, and 14 for corresponding visualizations in Appendix A. Finally, saliency map algorithms like Grad-CAM and Grad-CAM++ also enjoy improved stability through Gaussian smoothing in some scenarios according to Table 1. See Figures 11, 12, and 13 for visualizations of Integrated-Grad, Grad-CAM, and Grad-CAM++ in Appendix A.
@@ 6 Conclusion
In this work, we analyzed the stability fidelity trade-off in the widely-used Smooth-Grad interpretation map. Our results indicate the impact of Gaussian smoothing on the stability of saliency maps to the stochasticity of the training setting. On the other hand, we show Gaussian smoothing could lead to a higher discrepancy with the original gradient map. Therefore, Gaussian smoothing needs to be utilized properly, depending on the target application’s aimed stability and fidelity scores. Performing an analytical study of the effects of Gaussian smoothing on the sanity checks in [1] is an interesting future direction to our work. While [1] numerically shows that Smooth-Grad passes the sanity checks, an analysis of the $\sigma$ parameter in Gaussian smoothing and Smooth-Grad’s performance in the sanity checks will be a relevant topic for future study.
@@ Appendix A More Results
Due to the space limit of the main text, we put more results here. Figure 6, 7 show the relationship of stability error versus sigma, and fidelity error versus sigma in CIFAR10, respectively. Figure 8 displays two ResNet50 checkpoints trained with different recipes. Figure 9, 10, 14 show our findings extend to the case where two neural networks use different architectures (and naturally the training recipes also differ). Figure 11, 12, 13 shows the stabilizing effect of Gaussian smoothing for Integrated-Grad, Grad-CAM, and Grad-CAM++, respectively. These visualizations, together with Table 1 in the main text, provide empirical evidence that Gaussian smoothing can generally stabilize saliency maps across various settings. Additionally, Figure 15 visualizes that Smooth-Grad sacrifices fidelity with increasing smoothing parameter $\sigma$.
Figure 6: Relationship between stability error of saliency maps and sigma, evaluated on the CIFAR-10 test set. The shaded region indicates one standard deviation from the mean value. Here noise level $x_{\text{max}}-x_{\text{min}}$ represents the largest value range for all pixels, following [31].
Figure 8: Visualization of saliency maps of two pre-trained ResNet50 checkpoints provided by TorchVision [24]. The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 9: Visualization of saliency maps of pre-trained ResNet50 and Swin-T checkpoints provided by TorchVision [24]. The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 10: Visualization of saliency maps differences of pre-trained ResNet50 and ConvNeXt-tiny checkpoints provided by TorchVision [24]. The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
@@
Figure 12: Visualization of Grad-CAM and smoothed version of Grad-CAM differences on ImageNet, where the two neural networks are different pre-trained ResNet50 checkpoints provided by TorchVision [24]. The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 13: Visualization of Grad-CAM++ and smoothed version of Grad-CAM++ differences on ImageNet, where the two neural networks are different pre-trained ResNet50 checkpoints provided by TorchVision [24]. The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 14: Visualization of Simple-Grad and Smooth-Grad differences on ImageNet, where the three neural networks are pre-trained ResNet50, Swin-T, and ConvNeXt-tiny checkpoints provided by TorchVision [24]. The first row is the input image. The second, third, and fourth rows are Simple-Grad results. The fifth, sixth, and seventh rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Simple-Grad
Integrated-Grad
Figure 15: Visualization of saliency map becoming less similar to the unsmoothed version as $\sigma$ increases. We plot for both Simple-Grad and Integrated-Grad.
@@ C.1 Training Details
For ImageNet, We train 80 epochs using the SGD optimizer with a momentum of 0.9, the StepLR scheduler with a step size of 30 and a gamma of 0.1 to adjust the learning rate, and weight decay of 0.001. We use RandomResizedCrop to ensure images have shape $224\times 224$, and normalize using the mean and standard deviation computed from ImageNet. All these configurations are default ones in the training script provided by Pytorch.  [^1] https://github.com/pytorch/examples/blob/main/imagenet/main.py For CIFAR10, we train 100 epochs using the SGD optimizer with a momentum of 0.9, cosine learning rate annealing starting from 0.1 down to 0, and weight decay of 0.001. We normalize images using the mean and standard deviation of $0.5$ in all 3 channels. The base models we use are provided by TorchVision [24]. We run experiments on 4 GeForce RTX 3090 GPUs, the training time for a 500000 training subset takes within 12 hours. Except for the training process, most of our experiments run within one hour.
For each specified number of training data points $N$, we conduct the experiments three times, using seeds 10007, 5678, and 12345. In each run, the seed is employed to randomly shuffle the list of all indices. We then select the first $N$ indices to train one neural network, and the subsequent $N$ indices to train another. In this way, we can compute standard deviations for Figure 6, 2, 7, 3. For the experiment in 16 which requires 10 splits, we similarly use seed 10007 to shuffle and choose each consecutive $N$ samples as one split.
@@ C.2 Implementation of Saliency Map Algorithms
We use the saliency library provided by PAIR  [^2] https://github.com/PAIR-code/saliency to compute SimpleGrad, IntegratedGrad, and their smoothed versions, as well as visualize those saliency maps. We use the pytorch-grad-cam library  [^3] https://github.com/jacobgil/pytorch-grad-cam [11] to compute Grad-CAM and Grad-CAM++ since the former library does not natively support them. We set the desired layer to be model.layer4[-1] for ResNet50 and model.norm for Swin-T, both are close to final layers and our experiments show good visualization results under these choices. According to this library, we visualize a CAM figure as an RGB image by overlaying the cam mask on the image as a heatmap.

### v2

@@ 4.3 Fidelity of Gaussian Smoothed Saliency Maps
We then illustrate that this stability improvement comes at the expense of fidelity. As $\sigma$ increases, the Gaussian-smoothed saliency map diverges further from the ground-truth map.
First, we formally define the fidelity error $\epsilon_{\mathrm{fidelity}}(\operatorname{Sal},\sigma)$ of the $\sigma$-smoothed saliency map algorithm $\operatorname{Sal}$ as the largest possible expected $\ell_{2}$-distance between the smoothed and unsmoothed saliency maps for any training set $S$, labeled data sample $x,y$, where the expectation is over the randomness of the algorithm. Specifically, fidelity error is defined as:
$\sup_{S,x,y}\mathbb{E}_{A}\Bigl{[}\Big{\|}$ $\mathbb{E}_{z\sim N(0,\sigma^{2}I)}\left[\operatorname{Sal}_{A(S)}(x+z,y)\right]$
$-\operatorname{Sal}_{A(S)}(x,y)\Big{\|}\Bigr{]}$ (11)
@@ 5 NUMERICAL EXPERIMENTS
Figure 2: Relationship between stability error of saliency maps and sigma, evaluated on the CIFAR10 test set. The shaded region indicates one standard deviation from the mean value. Here noise level $x_{\text{max}}-x_{\text{min}}$ represents the largest value range for all pixels, following the practice of  Smilkov et al., (2017).
Figure 3: Relationship between stability error of saliency maps and sigma, evaluated on the ImageNet test set. The shaded region indicates one standard deviation from the mean value.
@@
The goal of our numerical experiments is to validate the stability-fidelity tradeoff incurred by choosing different smoothing factors $\sigma$. We evaluated different neural network architectures and different training set sizes on CIFAR10 (Krizhevsky and Hinton,, 2009) and ImageNet (Deng et al.,, 2009) datasets. We trained ResNet34 (He et al.,, 2015) for CIFAR10 and ResNet50 for ImageNet, due to different dataset scales. More detailed experiment configurations are in Appendix C.
@@ 5.1 Saliency Map Stability and Fidelity Numerical Results
$\mathbb{E}_{(x,y)\sim D}\mathbb{E}_{i=1}^{2}\bigl{\|}$ $\mathbb{E}_{z\sim N(0,\sigma^{2}I)}[\operatorname{Sal}_{A(S_{i})}(x+z,y)]$
$-\operatorname{Sal}_{A(S_{i})}(x,y)\bigr{\|}$ (13)
@@ 5.2 Visualizations of Saliency Map Stability and Fidelity
Figure 8: Visualization of Simple-Grad becoming less similar to the original unsmoothed map with the Gaussian smoothing as the factor $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})$ increases.
Additionally, we visualize the trend that a larger $\sigma$ results in a smoothed saliency map that is less similar to the vanilla version, as demonstrated in Figure 8 for Simple-Grad and in Figure 9 for Integrated-Grad.
The visualization results, along with our theoretical findings, indicate that Gaussian smoothing enhances the stability of saliency maps at the expense of fidelity. Therefore, the choice of $\sigma$ in practical applications should be carefully determined by considering this trade-off, since applications of saliency maps could require different levels of the two factors.
@@ 5.3 Empirical Results on Generalized Settings
Beyond the randomness of training settings, we further examine other sources that could cause instability in interpretation: different training recipes and model architectures. Extensive experiments across various settings, as shown in Table 1, empirically indicate that the stabilizing effect of Gaussian smoothing also holds in different scenarios. We use the structural similarity index measure SSIM (Wang et al.,, 2004) to evaluate the similarity of two saliency maps and a larger SSIM value indicates that the two maps are more perceptually similar. We additionally use top-k mIoU to evaluate the similarity by calculating mIoU between the most salient $k=500$ pixels of two maps.
In the top four rows, we train two ResNet50 models on two disjoint ImageNet training subsets of 500,000 samples each. For all other rows, we use pre-trained checkpoints provided by TorchVision (maintainers and contributors,, 2016). Since TorchVision provides two ResNet50 checkpoints trained with different schemes, we use this pair for the middle four rows of our table. For the ResNet50 models in the bottom rows, we use the checkpoint version 1.
First, when networks are trained on different training sets, Gaussian smoothing makes the resulting saliency maps more similar for both Simple-Grad and Integrated-Grad. For advanced saliency map algorithms such as Grad-CAM (Selvaraju et al.,, 2017) and Grad-CAM++ (Chattopadhay et al.,, 2018), the vanilla maps have high SSIM values due to the concentrated and block-like patterns, yet their top-k mIoU are still modest. Applying Gaussian smoothing occasionally causes significant visual changes in one saliency map but not the other, resulting in a slightly lower SSIM value for Grad-CAM. Second, the instability caused by different training recipes can be largely mitigated by Gaussian smoothing, with better SSIM and top-k mIoU scores. See the corresponding visualization for Simple-Grad in Figure 10 in Appendix A.
Third, the effect of smoothing holds for most cross-architecture experiments. Besides ResNet, we also evaluated more recent architectures such as Swin Transformer (Liu et al.,, 2021) and ConvNeXt (Liu et al.,, 2022). See Figures 11, 12, and 16 for corresponding visualizations in Appendix A. Finally, saliency map algorithms like Grad-CAM and Grad-CAM++ also enjoy improved stability through Gaussian smoothing in some scenarios according to Table 1. See Figures 13, 14, and 15 for visualizations of Integrated-Grad, Grad-CAM, and Grad-CAM++ in Appendix A. Figure 6 shows that the highest stabilizing effect of Gaussian smoothing can be achieved with a proper choice of $\sigma$.
@@ 6 CONCLUSION
In this work, we analyzed the stability fidelity trade-off in the widely-used Smooth-Grad interpretation map. Our results indicate the impact of Gaussian smoothing on the stability of saliency maps to the stochasticity of the training setting. On the other hand, we show Gaussian smoothing could lead to a higher discrepancy with the original gradient map. Therefore, Gaussian smoothing needs to be utilized properly, depending on the target application’s aimed stability and fidelity scores. Performing an analytical study of the effects of Gaussian smoothing on the sanity checks in Adebayo et al., (2018) is an interesting future direction to our work. While Adebayo et al., (2018) numerically shows that Smooth-Grad passes the sanity checks, an analysis of the $\sigma$ parameter in Gaussian smoothing and Smooth-Grad’s performance in the sanity checks will be a relevant topic for future study.
Acknowledgments
The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. Also, the authors would like to thank the anonymous reviewers for their constructive feedback and suggestions.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A More Results
Due to the space limit of the main text, we put more results here. Figure 9 visualizes the trend of less faithfulness to the original saliency map with increasing $\sigma$. Figure 10 displays two ResNet50 checkpoints trained with different recipes. Figure 11, 12, 16 show our findings extend to the case where two neural networks use different architectures (and naturally the training recipes also differ). Figure 13, 14, 15 shows the stabilizing effect of Gaussian smoothing for Integrated-Grad, Grad-CAM, and Grad-CAM++, respectively. These visualizations, together with Table 1 in the main text, provide empirical evidence that Gaussian smoothing can generally stabilize saliency maps across various settings.
Figure 9: Visualization of Integrated-Grad becoming less similar to the original unsmoothed map with the Gaussian smoothing as the factor $\sigma$ increases.
Figure 10: Visualization of saliency maps of two pre-trained ResNet50 checkpoints provided by TorchVision (maintainers and contributors,, 2016). The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 11: Visualization of saliency maps of pre-trained ResNet50 and Swin-T checkpoints provided by TorchVision (maintainers and contributors,, 2016). The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 12: Visualization of saliency maps differences of pre-trained ResNet50 and ConvNeXt-tiny checkpoints provided by TorchVision (maintainers and contributors,, 2016). The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
@@
Figure 14: Visualization of Grad-CAM and smoothed version of Grad-CAM differences on ImageNet, where the two neural networks are different pre-trained ResNet50 checkpoints provided by TorchVision (maintainers and contributors,, 2016). The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 15: Visualization of Grad-CAM++ and smoothed version of Grad-CAM++ differences on ImageNet, where the two neural networks are different pre-trained ResNet50 checkpoints provided by TorchVision (maintainers and contributors,, 2016). The first row is the input image. The second and third rows are Simple-Grad results. The fourth and fifth rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
Figure 16: Visualization of Simple-Grad and Smooth-Grad differences on ImageNet, where the three neural networks are pre-trained ResNet50, Swin-T, and ConvNeXt-tiny checkpoints provided by TorchVision (maintainers and contributors,, 2016). The first row is the input image. The second, third, and fourth rows are Simple-Grad results. The fifth, sixth, and seventh rows are Smooth-Grad with $\sigma/(x_{\mathrm{max}}-x_{\mathrm{min}})=0.15$.
In addition to Table 1 in the main text, we consider two more saliency map algorithms Gradient $\odot$ Input (Shrikumar et al.,, 2017) and MoreauGrad (Zhang and Farnia,, 2023) in Table 2 and can see the stabilizing effect of Gaussian smoothing still holds.
Table 2: Average SSIM and top-k mIoU scores for saliency maps from two neural networks evaluated on a random subset of the ImageNet test set. The two ResNet50 networks are trained on two different splits of the training set. Smoothed scores are computed after applying Gaussian smoothing to both saliency maps.
Algorithm SSIM Smoothed SSIM top-k mIoU Smoothed top-k mIoU
Gradient $\odot$ Input 0.264 0.323 0.092 0.154
MoreauGrad 0.621 0.785 0.118 0.156
@@ C.1 Training Details
For ImageNet, We train 80 epochs using the SGD optimizer with a momentum of 0.9, the StepLR scheduler with a step size of 30 and a gamma of 0.1 to adjust the learning rate, and weight decay of 0.001. We use RandomResizedCrop to ensure images have shape $224\times 224$, and normalize using the mean and standard deviation computed from ImageNet. All these configurations are default ones in the training script provided by Pytorch.  [^2] https://github.com/pytorch/examples/blob/main/imagenet/main.py For CIFAR10, we train 100 epochs using the SGD optimizer with a momentum of 0.9, cosine learning rate annealing starting from 0.1 down to 0, and weight decay of 0.001. We normalize images using the mean and standard deviation of $0.5$ in all 3 channels. The base models we use are provided by TorchVision (maintainers and contributors,, 2016). We run experiments on 4 GeForce RTX 3090 GPUs, the training time for a 500000 training subset takes within 12 hours. Except for the training process, most of our experiments run within one hour.
For each specified number of training data points $N$, we conduct the experiments three times, using seeds 10007, 5678, and 12345. In each run, the seed is employed to randomly shuffle the list of all indices. We then select the first $N$ indices to train one neural network, and the subsequent $N$ indices to train another. In this way, we can compute standard deviations for Figure 2, 3, 4, 5. For the experiment in 17 which requires 10 splits, we similarly use seed 10007 to shuffle and choose each consecutive $N$ samples as one split.
@@ C.2 Implementation of Saliency Map Algorithms
We use the saliency library provided by PAIR  [^3] https://github.com/PAIR-code/saliency to compute SimpleGrad, IntegratedGrad, and their smoothed versions, as well as visualize those saliency maps. We use the pytorch-grad-cam library  [^4] https://github.com/jacobgil/pytorch-grad-cam (Gildenblat and contributors,, 2021) to compute Grad-CAM and Grad-CAM++ since the former library does not natively support them. We set the desired layer to be model.layer4[-1] for ResNet50 and model.norm for Swin-T, both are close to final layers and our experiments show good visualization results under these choices. According to this library, we visualize a CAM figure as an RGB image by overlaying the cam mask on the image as a heatmap.